x-5y=-18 and x-7y=6
Answer by mananth(16949)

x-5y=-18 and x-7y=6
subtract the equations
x-5y-x+7y=-18-6
+2y=-24
y=-12
..
x-5y=-18
plug the value of y
x+60=-18
x=-78